# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 3.2 Causal consistency loss
It is not always possible (nor desirable) to achieve an exact transformation that guarantees consistency of low- and high-level models for almost all interventions. As a consequence, we allow for the consistency between models to be approximate. To ensure that this approximation is as accurate as possible, we minimize the expected KL divergence between the pushforward by the transformation $\tau$ of the low-level interventional distributions that we denote $\widehat{P}_{\tau}^{(\mathbf{i})}(Y,\mathbf{Z})=\tau_{\#}[P_{\mathcal{L}}^{(\mathbf{i})}(\mathbf{X})]$, and the corresponding interventional distribution of the high-level model $P^{(\omega(\mathbf{i}))}$. The associated consistency loss to be minimized is
$\mathcal{L}_{cons}=\mathbb{E}_{\mathbf{i}\sim P(\mathbf{i})}\left[KL\left(\widehat{P}_{\tau}^{(\mathbf{i})}(Y,\mathbf{Z})\|P^{(\omega(\mathbf{i}))}(Y,{\mathbf{Z}})\right)\right]\,,$ (4)
and has the following properties:
@@
The consistency loss is positive, invariant to invertible reparametrizations, and vanishes if and only if the transformation is exact for almost all interventions. It admits the following decomposition:
$\mathcal{L}_{cons}\!=\!\mathbb{E}_{i\sim P(\mathbf{i})}\!\!\left[KL\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)\right.\\ \!\left.+\mathbb{E}_{\mathbf{z}\sim\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\!\!\left[KL\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left({Y}|\mathbf{z})\right)||{P}^{(0)}\left({Y}|\mathbf{z}\right)\right)\right]\right]\,,$ (5)
@@
$\mathcal{L}_{rel}\!=\!\mathbb{E}_{\mathbf{i}\sim P(\mathbf{i})}\!\left[KL\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)\right]\!\leq\!\mathcal{L}_{cons}\,.$ (6)
The property of reparametrization invariance, detailed in the proof in App. B, refers to transformations of the pairs $(\tau,f_{\bm{\gamma}})$ that leave the composition $f_{\bm{\gamma}}\circ\tau$ invariant. In the $n=1$ linear setting (see Sec. 3.3), this corresponds to invariance by multiplicative rescaling.
@@
Lastly, the causal relevance loss $\mathcal{L}_{rel}$ assesses whether the variations of the target $Y$ due to low-level interventions are well captured by high-level interventions, on average over the prior $P(\mathbf{i})$. Its upper bound by $\mathcal{L}_{cons}$ ensures that by optimizing for consistency, we also indirectly promote effective “explanation” of the variations in the target density resulting from low-level interventions. We can thus choose $P(\mathbf{i})$ to make the most relevant interventions more likely according to domain knowledge, such that optimizing the loss will steer towards a solution capturing the most domain-relevant variations of the target.
@@ 3.3 Linear reduction with shift interventions
We further constrain the setting to be able to study the solution minimizing $\mathcal{L}_{cons}$ analytically and get insights into how TCR works. In this section, we describe the approach for the 1D TCR, with $n=1$ high-level cause. The case $n>1$ is discussed in App. C.
@@
We use boldface, e.g. $\mathbf{i}$, for column vectors, and $\mathbf{i}_{S}$ for the subvector of $\mathbf{i}$ restricted to the components in set $S$. When a vector, say $\bm{\tau}_{k}$, is associated to high-level SCM component $k$ of a constructive transformation with alignment $\pi$, $\bar{\bm{\tau}}_{k}$ indicates the restriction of $\bm{\tau}_{k}$ to components in $\pi(k)$.
@@
$\mathbf{X}\mapsto\begin{bmatrix}Y\\ {\mathbf{Z}}\end{bmatrix}=\begin{bmatrix}\bm{\tau}_{0}^{\top}\\ \bm{\tau}_{1}^{\top}\end{bmatrix}\mathbf{X}=\begin{bmatrix}\bar{\bm{\tau}}_{0}^{\top}\mathbf{X}_{\pi(0)}\,,&\bar{\bm{\tau}}_{1}^{\top}\mathbf{X}_{\pi(1)}\end{bmatrix}^{\top}.$
@@
We focus on shift interventions and map the vector $\bm{i}$ of low-level interventions on the nodes in $\pi(1)$ to a scalar shift intervention on the mechanism of $Z_{1}$. We assume this map $\omega_{1}$ to be linear and thus define the vector ${\bm{\omega}}_{1}$ such that the map is given by
$\omega_{1}(\mathbf{i})={\bm{\omega}}_{1}^{\top}\bm{i}=\bar{{\bm{\omega}}}_{1}^{\top}\bm{i}_{\pi(1)}\,.$
Because it is a root node, intervening amounts to shifting the marginal distribution from $P^{(0)}(Z_{1})$ to $P^{(\omega_{1}(\mathbf{i}))}(Z_{1})=P^{(0)}(Z_{1}-\omega_{1}(\mathbf{i}))$.
@@
In order to form a constructive transformation we need to impose some restrictions on the alignment: (1) $\pi(0)$ must include the support of $\bm{\tau}_{0}$, which is given; (2) $\pi(1)$ must not overlap with $\pi(0)$. Beyond that, there are potential degrees of freedom for $\pi$, and users may want to incorporate domain knowledge as well as interpretability constraints to reduce the variables included in $\pi(1)$. This can be done by fixing $\pi$ prior to learning, or by adapting it during learning, using for example additional terms to the loss enforcing sparsity. Such sparsity enforcing regularizers are particularly useful for the case of $n\geq 2$ high-level causes, as shown in App. C. For the next identifiability results of Sec. 3.4, we will first fix $\pi$ and show that different choices may lead to indeterminacies of the TCR solution.
@@
The solutions minimizing the loss of Eq. (4), may depend on the choice of the prior $P(\mathbf{i})$, and in particular on which variables are actually intervened on. For our theoretical analysis, we assume the prior has the following form.
@@
There exists a (unique) subset ${\Omega}$ of intervened low-level variables such that $P(\mathbf{i}_{{\Omega}})$ has a density with respect to the Lebesgue measure, with support covering a neighborhood of zero (i.e. the unintervened case). Components of $\mathbf{i}$ associated to unintervened variables (outside of $\Omega$) take the deterministic value $i=0$.
@@ 3.4 Identifiability results
If we assume the low-level model is linear Gaussian of the form $\mathbf{X}_{\pi(1)}\to\mathbf{X}_{\pi(0)}$, we can show the existence and uniqueness of the solution, under the condition that $\pi(1)$ contains all and only intervened variables.
Figure 2: TCR solutions on a chain graph. Arrows indicate the non-zero coefficients of each map. (a) Unique solution $\bm{\tau}_{1}$ when interventions are on all nodes but the target. (b) Two solutions $\bm{\tau}_{1}$ and $\bm{\tau}_{1}^{\prime}$ when interventions are on the first two nodes only.
@@
$\mathbf{X}\coloneqq A\mathbf{X}+\mathbf{U}+\mathbf{i}\,,\quad U_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})\,,\,\,\mathbf{i}\sim P(\mathbf{i})$
@@
Under Assum. 3.2, if $\Omega=\pi(1)$, and for a given choice of target $Y=\tau_{0}^{\top}\mathbf{X}=\bar{\tau}_{0}^{\top}\mathbf{X}_{\pi(0)}$, there is a unique linear 1D TCR, up to a multiplicative constant satisfying $\mathcal{L}_{cons}=0$. It is given by
$\bar{\bm{\tau}}_{1}=$ $\,A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}$ (7)
@@
Additionally, Assum. 3.2 can be replaced by Assum. B.2, which requires only a finite set of $\pi(1)$ interventions.
Interpretation.
This solution is illustrated in Fig. 2(a) for a chain graph with $\pi(0)=\{4\}$ and $\pi(1)=\{1,2,3\}$. We can see that $\bm{\tau}_{1}$ puts all its weight on the unique parent of $X_{4}=Y$. This can be seen as enforcing the mechanism consistency part of $\mathcal{L}_{cons}$. In contrast, ${\bm{\omega}}_{1}$ puts weight on all intervened variables, to enforce cause consistency of changes in $Z_{1}$ due to all interventions. In general, ${\bm{\omega}}_{1}$ and $\bm{\tau}_{1}$ will differ whenever there are causal interactions between variables in $\pi(1)$, encoded by $A_{11}$. See App. D for more details. This example shows that the set of intervened variables that affect shifts of $Z_{1}$ through ${\bm{\omega}}_{1}$ may strongly differ from those mapped by the reduction $\bm{\tau}_{1}$. This may lead to conflicting interpretations of $Z_{1}$, and will be addressed by regularization in Sec. 4.
Our next result shows that we loose identifiability of the TCR if we drop the assumption that all and only variables in $\pi(1)$ are intervened on.
Proposition 3.4.
Consider the setting of Prop. 3.3 with the exception that $\Omega\subsetneqq\pi(1)$ such that there is now a non-empty subset $S=\pi(1)\setminus\Omega$, such that $\mathbf{X}_{\Omega}\to\mathbf{X}_{S}\to\mathbf{X}_{\pi(0)}$. Then there exist at least two different linear 1D TCR such that $\mathcal{L}_{cons}=0$.
This result can also be illustrated with a chain graph, as shown in Fig. 2(b). If the parent node $X_{3}$ of $Y=X_{4}$ is unintervened, then one may choose either $Z_{1}=X_{2}$ or $Z_{1}=X_{3}$ (matching the solution of Fig. 2(a)) to minimize $\mathcal{L}_{cons}$. This is because both variables are equivalently mediating all performed interventions to $X_{4}$. Note that each choice has it own benefit: $Z_{1}=X_{3}$, as a direct parent of $Y$, is a better statistical predictor of the value of $Y$. However, if we focus on causal interpretability of the high-level representation, $Z_{1}=X_{2}$ is preferable because it is one of the variables intervened on at the low-level as enforced by the prior $P(\mathbf{i})$, and such that it will be associated to a non-zero weight in ${\bm{\omega}}_{1}$ for any solution satisfying $\mathcal{L}_{cons}=0$.
This demonstrates that the choice of $\pi(1)$ may introduce ambiguities when searching for a consistent solution, and selecting where and how to intervene in the system plays a crucial role. While it is possible to restrict $\pi(1)$ to nodes that are directly intervened upon, some nodes may in practice be intervened on sensu stricto, while having a minimal impact on $Y$, yet they may carry most of the weight of $\bm{\tau}_{1}$. This again suggests a potential “inhomogenity” between the high-level representation of the cause, with ${\bm{\omega}}_{1}$ and $\bm{\tau}_{1}$ reflecting different aspects of the low-level causes. To address this, we can enforce homogeneity of the solution by requiring ${\bm{\omega}}$ and $\bm{\tau}$ to be as similar as possible, which we implement as an optional regularization term of our learning objective in Sec. 4.
@@ 4 LINEAR TCR ALGORITHM
In this section, we introduce an algorithm to learn a linear reduction with shift interventions.
Gaussian approximation of consistency loss.
Since the KL divergence is challenging to learn non-parametrically, we make a Gaussian assumption on the densities. This allows us to obtain an analytic expression for the loss based on second order statistics (see expression in App. E.1). In order to promote interpretability, we supplement $\mathcal{L}_{cons}$ with a homogeneity regularizer.
Interventional homogeneity loss.
Importantly, $\omega$ and $\tau$ can assign strong coefficients on different sets of variables. This is related to the causal structure of the set of variables constituting $Z_{k}$. This poses interpretability challenges, as the variables acted on are different from the variables used for prediction. To align those sets of variables, we introduce the homogeneity loss that is maximal when $\omega=\tau$, and scale invariant:
${\mathcal{L}_{hom}=\sum_{k}\left\langle\frac{{\bm{\omega}}_{k}}{\|{\bm{\omega}}_{k}\|},\frac{\bm{\tau}_{k}}{\|\bm{\tau}_{k}\|}\right\rangle.}$ (9)
Gathering the two losses, we get the total objective
$\underset{\bm{\gamma},\bm{\tau}_{1},{\bm{\omega}}_{1}}{\mbox{minimize }}\mathcal{L}_{tot}=\mathcal{L}_{cons}-\eta_{h}\mathcal{L}_{hom}\,.$
The learning procedure is described in Algorithm 1.
Algorithm 1 Linear TCR (LCPR)
Input $\lambda$: learning rate, $P(\mathbf{i})$: intervention prior, Simulate$(\bm{\theta},\mathbf{i},n_{sim})$: function returning $n_{sim}$ paths, $N_{ite}$: number of iterations, $B$: simulation paths batch size, $B_{\mathbf{i}}$: intervention batch size.
Output Estimated parameters $(\bm{\tau}_{1},{\bm{\omega}}_{1},\bm{\gamma})$.
@@
      $X_{l}=(\mathbf{x}^{1},..,\mathbf{x}^{B}),Y_{l}\leftarrow Simulate(\bm{\theta},\mathbf{i}_{l},B)$
      $X\leftarrow[X[:],X_{l}]$
      $Y\leftarrow[Y[:],Y_{l}]$
      $I=[I[:],\mathbf{i}_{l}]$
   $L_{tot}\leftarrow ComputeLoss(X,Y,I,\bm{\tau}_{1},\omega_{1},\bm{\gamma})$
   $\nabla_{\bm{\gamma}},\nabla_{\bm{\tau}}\leftarrow ComputeLossGradient(L_{tot})$

### v2

@@ 3.2 Causal consistency loss
It is not always possible to achieve an exact transformation that guarantees consistency of low- and high-level models for almost all interventions. As a consequence, we allow for the consistency between models to be approximate. To ensure that this approximation is as accurate as possible, we minimize the expected KL divergence between the pushforward by the transformation $\tau$ of the low-level interventional distributions that we denote $\widehat{P}_{\tau}^{(\mathbf{i})}(Y,\mathbf{Z})=\tau_{\#}[P_{\mathcal{L}}^{(\mathbf{i})}(\mathbf{X})]$, and the corresponding interventional distribution of the high-level model $P^{(\omega(\mathbf{i}))}$, leading to the consistency loss
$\mathcal{L}_{\mathrm{cons}}=\mathbb{E}_{\mathbf{i}\sim P(\mathbf{i})}\left[\mathrm{KL}\left(\widehat{P}_{\tau}^{(\mathbf{i})}(Y,\mathbf{Z})\|P^{(\omega(\mathbf{i}))}(Y,{\mathbf{Z}})\right)\right]\,.$ (4)
Other losses have been previously suggested to enforce consistency. Beckers et al. (2020) propose to take a maximum over interventions, whereas we take the expectation in our loss, thus focusing the CMR on the average performance rather than the worst case. Rischel and Weichwald (2021) and Zennaro et al. (2023) use the Jensen-Shannon (JS) divergence in the context of finite models. Instead, we choose the KL divergence because, contrary to JS, it leads to a tractable expression under Gaussian assumptions. Moreover, the proposed consistency loss (4) has the following properties.
@@
The consistency loss is positive, invariant to invertible reparametrizations (see Def. D.1), and vanishes if and only if the transformation is exact for almost all interventions. It decomposes as
$\mathcal{L}_{\mathrm{cons}}\!=\!\mathbb{E}_{i\sim P(\mathbf{i})}\!\!\left[\mathrm{KL}\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)\right.\\ \!\left.+\mathbb{E}_{\mathbf{z}\sim\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\!\!\left[\mathrm{KL}\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left({Y}|\mathbf{z})\right)||{P}^{(0)}\left({Y}|\mathbf{z}\right)\right)\right]\right]\,,$ (5)
@@
$\mathcal{L}_{\mathrm{rel}}\!=\!\mathbb{E}_{\mathbf{i}\sim P(\mathbf{i})}\!\left[\mathrm{KL}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)\right]\!\!\leq\!\!\mathcal{L}_{\mathrm{cons}}\,.$ (6)
Reparametrization invariance (see Def. D.1) refers to transformations of the pairs $(\tau,f_{\bm{\gamma}})$ that leave the composition $f_{\bm{\gamma}}\circ\tau$ invariant. In the $n=1$ linear setting (see Sec. 3.3), this corresponds to invariance by multiplicative rescaling. This guarantees that equivalent high-level causal descriptions are treated equally by the loss.
@@
Lastly, the causal relevance loss $\mathcal{L}_{rel}$ assesses whether the variations of the target $Y$ due to low-level interventions are well-captured by high-level interventions, on average over the prior $P(\mathbf{i})$. Its upper bounding by $\mathcal{L}_{\mathrm{cons}}$ ensures that by optimizing for consistency, we also indirectly promote effective “explanations” of the variations in the target density resulting from low-level interventions. We can thus choose $P(\mathbf{i})$ to make the most relevant interventions more likely according to domain knowledge, such that optimizing the loss will steer towards a solution capturing the most domain-relevant variations of the target.
@@ 3.3 Linear reduction with shift interventions
We further constrain the setting to be able to study the solution minimizing $\mathcal{L}_{\mathrm{cons}}$ analytically and get insights into the properties of TCR.
@@
When a vector, say $\bm{\tau}_{k}$, is associated to a high-level SCM component $k$ of a constructive transformation with alignment $\pi$, $\bar{\bm{\tau}}_{k}$ indicates the restriction of $\bm{\tau}_{k}$ to components in $\pi(k)$. The number of elements in a set $S$ is #$\#S$.
@@
$\mathbf{X}\mapsto\begin{bmatrix}Y\\ {\mathbf{Z}}\end{bmatrix}=\begin{bmatrix}\bm{\tau}_{0}^{\top}\\ \vdots,\\ \bm{\tau}_{n}^{\top}\end{bmatrix}\mathbf{X}=\begin{bmatrix}\bar{\bm{\tau}}_{0}^{\top}\mathbf{X}_{\pi(0)}\,,&\dots\bar{\bm{\tau}}_{n}^{\top}\mathbf{X}_{\pi(n)}\end{bmatrix}^{\top}.$
@@
We focus on shift interventions and map the vector $\bm{i}$ of low-level interventions on the nodes in $\pi(k)$ to a scalar shift intervention on the mechanism of each $Z_{k}$. We assume each map $\omega_{k}$ to be linear with vector ${\bm{\omega}}_{k}$ such that
$\omega_{k}(\mathbf{i})={\bm{\omega}}_{k}^{\top}\bm{i}=\bar{{\bm{\omega}}}_{k}^{\top}\bm{i}_{\pi(k)}\,.$
Because high-level causes are root nodes, intervening amounts to shifting the marginal distribution from $P^{(0)}(Z_{k})$ to $P^{(\omega_{k}(\mathbf{i}))}(Z_{k})=P^{(0)}(Z_{k}-\omega_{k}(\mathbf{i}))$.
@@
There are potential degrees of freedom for $\pi$, and users may want to incorporate domain knowledge as well as interpretability constraints to reduce the variables included in $\cup_{k\neq 0}\pi(k)$. In practice, we learn the distribution of the low-level variables among the $\pi(k)$ using regularization (see Sec. 4).
High-level mechanism.
We use an interpretable affine high-level causal mechanism $f_{\gamma}$, such that
$Y\coloneqq\sum_{k}\alpha_{k}Z_{k}+R_{0}+\beta\,,\quad\alpha_{1},\dots,\alpha_{n},\beta\in\mathbb{R}\,.$ (7)
@@
The solutions minimizing the loss of Eq. (4), may depend on the choice of the prior $P(\mathbf{i})$, and in particular on which variables are actually intervened on. Let ${\Omega}$ denote the subset of indices of low-level variables that are intervened on with non-zero probability. The components of $\mathbf{i}$ whose index does not belong to $\Omega$ thus take value $i=0$ with probability one. We provide identifiability guaranties under two kinds of assumptions.
@@
$P(\mathbf{i}_{{\Omega}})$ has a density with respect to the Lebesgue measure, with support covering a neighborhood of zero (i.e. the unintervened case).
Assumption 3.3.
The unintervened setting $\mathbf{i}_{\Omega}=\bm{0}$ occurs with non-zero probability. Additionally there are at least #$\#\Omega$ distinct interventions happening with non-zero probability, corresponding to a family of values of the vector of $\mathbf{i}_{\Omega}$ with full rank #$\#\Omega$.
While Assum. 3.2 depicts a practical setting where interventions are drawn from prior densities that reflect the prior knowledge on how likely those are, Assum. 3.3 allows addressing a classical question in causal representation learning: How many distinct interventions are needed to learn the representation?
@@ 3.4 Identifiability results
If we assume the low-level model is linear Gaussian of the form $\mathbf{X}_{\Omega}\to\mathbf{X}_{\pi(0)}$, we can show the existence and uniqueness of the solution.
@@
$\mathbf{X}\coloneqq A\mathbf{X}+\mathbf{U}+\mathbf{i}\,,\,\,\mathbf{i}\sim P(\mathbf{i})\,,\quad U_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})\,,\sigma_{k}>0\,,$
@@
$\mathbf{X}=\begin{bmatrix}\mathbf{X}_{\pi(0)}\\ \mathbf{X}_{\Omega}\end{bmatrix}\,,\quad A=\left[\begin{matrix}A_{00}\,&A_{0\Omega}\\ \bm{0}\,&A_{\Omega\Omega}\end{matrix}\right]\,.$
Given an arbitrary choice of linear scalar target of the form $Y=\tau_{0}^{\top}\mathbf{X}=\bar{\tau}_{0}^{\top}\mathbf{X}_{\pi(0)}$ and under Assum. 3.2 or Assum. 3.3, there is a unique linear 1-cause TCR (up to a multiplicative constant) satisfying $\mathcal{L}_{\mathrm{cons}}=0$. It is given by
$\pi(1)=$ $\,\Omega\,,$ (8)
$\bar{\bm{\tau}}_{1}=$ $\,A_{0\Omega}^{\top}(I_{\#\pi(0)}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}\,,$ (9)
$\text{and}\quad\bar{{\bm{\omega}}}_{1}=$ $\,(I_{\#\Omega}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\,.$ (10)
Moreover, let $n_{max}$ be the maximum number $n$ such that a linear $n$-cause TCR can achieve $\mathcal{L}_{\mathrm{cons}}=0$. If there are no cancellations [^2] Causal pathways cancel if the linear coefficients quantifying the influence of a node on $Y$ along different directed paths of the low-level SCM sum to zero. Assuming no cancellations is akin to assuming no faithfulness violations and generically satisfied (Sprites et al., 2001, Theorem 3.2). among causal pathways from each node in $\mbox{supp}(\bar{{\bm{\omega}}}_{1})$ of Eq. (10) towards $Y$, then the $n_{max}$-cause TCR is unique up to rescaling and permutation of the causes.
This result provides guaranties for having a unique ground-truth solution in case exact transformations can be achieved. The main assumption is the absence of feedback influences from the target set $\pi(0)$ to candidate causes. However, cycles and confounding are allowed in the low-level model, contrary to the learned high-level model. The 1-cause solution is easiest to obtain. The study of simple SCMs (App. D.2 and App. D.3) provides some insights on the form of the analytical solution. Additional results show that we lose identifiability of the TCR if we drop the assumption that not all variables in $\pi(1)$ are intervened on (see App. D.4). The $n$-cause solution is essentially a partition of the 1-cause solution that enforces independence between them.
Under the same model assumptions, the resulting constructive transformation can be associated with a constructive causal abstraction, as shown in Proposition D.3. This corresponds to a particular case of low soft abstraction introduced by Massidda et al. (2023, Def. 9).
Example 3.5 (Linear chain).
To illustrate the solutions in Prop. 3.4, we consider a linear chain
$\underbrace{X_{1}\rightarrow X_{2}\rightarrow X_{3}}_{X_{\Omega}}\rightarrow\underbrace{X_{4}}_{X_{\pi(0)}}$
@@
$\bar{\bm{\tau}}_{1}=\begin{pmatrix}0&0&1\end{pmatrix}^{\top}\quad\text{and}\quad\bar{{\bm{\omega}}}_{1}=\begin{pmatrix}1&1&1\end{pmatrix}^{\top}\,.$ (11)
$\bar{\bm{\tau}}_{0}$ puts all its weight on the direct parent of target $X_{4}$ because it mediates all causal influences. In contrast, $\bar{{\bm{\omega}}}_{1}$ puts weight on all variables in $\Omega$ because interventions on any of them influence $X_{4}$
@@ 4 LINEAR TCR ALGORITHM
In this section, we introduce an algorithm to learn a linear targeted causal reduction with shift interventions.
Algorithm 1 Linear TCR (LTCR)
Input $\lambda$: learning rate, $P(\mathbf{i})$: intervention prior, Simulate$(\bm{\theta},\mathbf{i},n_{\mathrm{sim}})$: function returning $n_{\mathrm{sim}}$ paths, $N_{\mathrm{ite}}$: No. epochs, $B$, $B_{\mathbf{i}}$: simulation/intervention batch size.
@@
      $X_{l}=(\mathbf{x}^{1},..,\mathbf{x}^{B})$; $Y_{l}\leftarrow Simulate(\bm{\theta},\mathbf{i}_{l},B)$
      $X\leftarrow[X[:],X_{l}]$; $Y\leftarrow[Y[:],Y_{l}]$; $I\leftarrow[I[:],\mathbf{i}_{l}]$
   $L_{\mathrm{tot}}\leftarrow ComputeLoss(X,Y,I,\bm{\tau}_{1},\omega_{1},\bm{\gamma})$
   $\nabla_{\bm{\gamma}},\nabla_{\bm{\tau}}\leftarrow ComputeLossGradient(L_{\mathrm{tot}})$
@@
Output Estimated parameters $(\bm{\tau}_{1},{\bm{\omega}}_{1},\bm{\gamma})$.
(a) Comparison between learned and analytical solutions 1-cause TCR Average cosine similarity to the analytical solutions over 20 runs. Each run corresponds to one draw of adjacency matrix parameters. The shaded areas show the range between the minimum and maximum values. The dashed gray line corresponds to perfect similarity.
(b) Two-branch linear model. Learned $\tau$- and $\omega$ parameters for a TCR with two high-level variables for a linear Gaussian low-level model with $N{=}10$. The solid lines show the parameters for $Z_{1}$ and the dashed lines those for $Z_{2}$. The parameters are averaged over 20 runs where each run corresponds to one draw of adjacency matrix parameters. The inlay shows the causal structure of the low-level model, where two groups of variables $G_{1}$ and $G_{2}$ form two independent chains causing the target $X_{10}{=}Y$.
Figure 2: Toy example experiments.
Gaussian approximation of consistency loss.
Since the KL divergence is challenging to compute in non-parametric settings, we make a Gaussian assumption on the densities. This allows us to obtain an analytic expression for the loss based on second order statistics (see expression in App. E.1).
Figure 3: Double well experiment. (a) Experimental setup with a ball moving in a double well potential subject to linear friction. (b) Pushforward density of the high-level cause for the two settings: one where no intervention is applied (unintervened), and the other with an applied shift intervention. (c, d) Learned parameters, $\tau$ and $\omega$, respectively. The learned high-level mechanism is $f(Z_{1})\approx 1.37Z_{1}+0.45$ (e) Samples in phase space (position vs. velocity) for the first 20 time points. The color indicates whether the high-level model predicts the ball to end up in the right (pink) or right well (turquoise). (f, g) Samples from the unintervened setting and the corresponding estimated density. (h) Estimated density for one intervened setting.
Overlap loss.
To ensure differentiability of the reduction maps we do not implement the alignment $\pi$ explicitly, but encourage non-overlapping reduction maps via the regularizer
$\mathcal{L}_{\mathrm{ov}}=\sum_{k<l}\left(\left\langle\frac{|\bm{\tau}_{k}|}{\|\bm{\tau}_{k}\|},\frac{|\bm{\tau}_{l}|}{\|\bm{\tau}_{l}\|}\right\rangle+\left\langle\frac{|{\bm{\omega}}_{k}|}{\|{\bm{\omega}}_{k}\|},\frac{|{\bm{\omega}}_{l}|}{\|{\bm{\omega}}_{l}\|}\right\rangle\right),$ (12)
where ||$\cdot$|| ⋅ | is the element-wise absolute value.
Balancing loss.
Minimizing the Gaussian approximation of the consistency loss together with overlap regularization (12) there is nothing preventing the solution from attributing all non-zero weights in the $\tau$- and $\omega$ maps to one high-level variable while ignoring all others. In order to prevent such a collapse, we minimize stark differences between the high-level variables through the balancing term
$\mathcal{L}_{\mathrm{bal}}=\Biggl{(}\ \frac{\sqrt{\sum_{k}\|\alpha_{k}\bm{\tau}_{k}\|_{2}^{2}}}{\sum_{k}\|\alpha_{k}\bm{\tau}_{k}\|_{2}}+\frac{\sqrt{\sum_{k}\|\alpha_{k}{\bm{\omega}}_{k}\|_{2}^{2}}}{\sum_{k}\|\alpha_{k}{\bm{\omega}}_{k}\|_{2}}\ \Biggr{)}\,,$ (13)
where $\alpha_{k}$ is the coefficient in the linear high-level mechanism corresponding to variable $Z_{k}$.
Gathering the losses, we get the total objective
$\underset{\bm{\gamma},\bm{\tau},{\bm{\omega}}}{\mbox{minimize }}\mathcal{L}_{\mathrm{tot}}=\mathcal{L}_{\mathrm{cons}}+\eta_{\mathrm{ov}}\mathcal{L}_{\mathrm{ov}}+\eta_{\mathrm{bal}}\mathcal{L}_{\mathrm{bal}}\,.$ (14)
The learning procedure is described in Algorithm 1.